# EMAGE: Towards Unified Holistic Co-Speech Gesture Generation via Expressive Masked Audio Gesture Modeling

### 02 Jan 2024 v2

@@ Appendix A Evaluation Metrics
where $\mu_{r}$ and $\Sigma_{r}$ represent the first and second moments of the latent features distribution $z_{r}$ of real human gestures $\mathbf{g}$, and $\mu_{g}$ and $\Sigma_{g}$ represent the first and second moment of the latent features distribution $z_{g}$ of generated gestures $\hat{\mathbf{g}}$. We use a Skeleton CNN (SKCNN) based encoder [1] and a Full CNN-based decoder as the autoencoder pretrained network. The network is pretrained on both BEATX-Standard and BEATX-Additional Data. The choice of SKCNN over a Full CNN encoder is due to its enhanced capability in capturing gesture features, as indicated by a lower reconstruction MSE loss (0.095 compared to 0.103).
@@
Appendix B BEATX Dataset Details
@@
The original BEAT dataset, as described by [40], contains 76 hours of data for 30 speakers. We exclude speakers 8, 14, 19, 23, and 29, which account for 16 hours of data, due to noise in the finger data, leaving 60 hours of data for 25 speakers (12 female and 13 male). The speech and conversation portions are categorized into BEATX-standard and BEATX-additional, containing 27 and 33 hours respectively. We adopt an 85%, 7.5%, and 7.5% split for BEATX-standard, maintaining the same ratio for each speaker. BEATX-additional is utilized for further improving the network’s robustness. The results presented in this paper are based on training with BEATX-standard only. The dataset includes 1762 sequences with an average length of 65.66 seconds per sequence. Each recording in a sequence is a continuous answer to a daily question. Additionally, we report a comparison between TalkShow [65] and our dataset in terms of Diversity and Beat Constancy (BC), as shown in Table 8.
@@ Appendix C Baselines Reproduction Details
Table 9: Vertex Errors (MSE) with Different Training Methods. ‘FF’ and ‘CT’ refer to FaceFormer [18] and CodeTalker [62], respectively. ‘TF’, ‘AR’, and ‘NonAR’ represent Teacher-Force, AutoRegressive, and Non-AutoRegressive training, respectively. We train on the VOCA dataset with a vertex loss, and BEATX with a FLAME parameter loss combined with a vertex loss. Results indicate that the same method performs differently when using the two representations; in BEATX, non-autoregressive training demonstrates superior performance. The average MSE is calculated on 5023 and 10475 vertices for VOCA and BEATX, respectively:
@@ Appendix E Visualization Blender Add-on
For straightforward visualization of our BEATX dataset, we utilize the SMPL-X Blender add-on [50]. As the latest SMPL-X add-on does not support the full range of facial expressions for SMPL-X, we extract 300 expression meshes from the original SMPL-X model and added them as individual blendshape targets into the SMPL-X model within the Blender add-on.
Appendix F Video, Data, and Codes
In the supplementary materials, we provide low-resolution videos, examples of the dataset, and codes. Due to its size, the complete version of the dataset is not included in the submission.

### 25 Mar 2024 v3

@@ Appendix A Evaluation Metrics
where $\mu_{r}$ and $\Sigma_{r}$ represent the first and second moments of the latent features distribution $z_{r}$ of real human gestures $\mathbf{g}$, and $\mu_{g}$ and $\Sigma_{g}$ represent the first and second moment of the latent features distribution $z_{g}$ of generated gestures $\hat{\mathbf{g}}$. We use a Skeleton CNN (SKCNN) based encoder [1] and a Full CNN-based decoder as the autoencoder pretrained network. The network is pretrained on both BEAT2-Standard and BEAT2-Additional Data. The choice of SKCNN over a Full CNN encoder is due to its enhanced capability in capturing gesture features, as indicated by a lower reconstruction MSE loss (0.095 compared to 0.103).
@@
Appendix B BEAT2 Dataset Details
@@
The original BEAT dataset, as described by [39], contains 76 hours of data for 30 speakers. We exclude speakers 8, 14, 19, 23, and 29, which account for 16 hours of data, due to noise in the finger data, leaving 60 hours of data for 25 speakers (12 female and 13 male). The speech and conversation portions are categorized into BEAT2-standard and BEAT2-additional, containing 27 and 33 hours respectively. We adopt an 85%, 7.5%, and 7.5% split for BEAT2-standard, maintaining the same ratio for each speaker. BEAT2-additional is utilized for further improving the network’s robustness. The results presented in this paper are based on training with BEAT2-standard speaker-2 only. The dataset includes 1762 sequences with an average length of 65.66 seconds per sequence. Each recording in a sequence is a continuous answer to a daily question. Additionally, we report a comparison between TalkShow [64] and our dataset in terms of Diversity and Beat Constancy (BC), as shown in Table 8.
@@ Appendix C Baselines Reproduction Details
Table 9: Vertex Errors (MSE) with Different Training Methods. ‘FF’ and ‘CT’ refer to FaceFormer [20] and CodeTalker [60], respectively. ‘TF’, ‘AR’, and ‘NonAR’ represent Teacher-Force, AutoRegressive, and Non-AutoRegressive training, respectively. We train on the VOCA dataset with a vertex loss, and BEAT2 with a FLAME parameter loss combined with a vertex loss. Results indicate that the same method performs differently when using the two representations; in BEAT2, non-autoregressive training demonstrates superior performance. The average MSE is calculated on 5023 and 10475 vertices for VOCA and BEAT2, respectively:
@@ Appendix E Visualization Blender Add-on
For straightforward visualization of our BEAT2 dataset, we utilize the SMPL-X Blender add-on [49]. As the latest SMPL-X add-on does not support the full range of facial expressions for SMPL-X, we extract 300 expression meshes from the original SMPL-X model and added them as individual blendshape targets into the SMPL-X model within the Blender add-on.
Appendix F Training time
We report the training time on a single L4, V100 and 4090 with a batch size (BS) of 64 for the best performance.
1-speaker 25-speaker
EMAGE 1-epoch 400-epoch 1-epoch 100-epoch Mem. BS
L4 (24G) 239s 26.5h 3197s 89.6h 20.1G 64
V100 (32G) 155s 17.2h 2073s 58.1h 20.1G 64
4090 (24G) 72s 8.0h 963s 27.1h 20.1G 64
Addtionally, pretraining of the 5 $\times$ VQVAEs for face, hands, upper body, lower body, and global motion would take 22.4 hours on 5 $\times$ 4090 GPUs.
1-speaker 25-speaker
VQVAEs $\times$ 1 1-epoch 700-epoch 1-epoch 100-epoch Mem. BS
L4 (24G) 200s 39.5h 2760s 74.4h 13.8G 64
V100 (32G) 131s 25.5h 1727s 48.0h 13.8G 64
4090 (24G) 61s 11.9h 802s 22.4h 13.8G 64
Appendix G Importance of lower body motion
Lower body motion allows gestures semantically aligned with the content of the audio to achieve more vivid and impressive results, e.g., “hiking in nature” with a walking gesture, “playing football” with a kicking motion; see figure below. Compared with the upper body, it is more weakly related to the audio, but it still has connections in the above cases.
In the implementation of EMAGE, we first obtain the latents of different body components with separate MLPs. Then, the lower body motion decoder leverages all latents of “audio”, “upper body”, and “hands” for cross-attention based lower-body motion decoding. We have also observed that decoding directly from audio increases diversity but reduces the coherence of the results on BEATv1.3.
[…]
audio only 6.209 6.683 13.714 1.183 8.788
audio + upper + hands 5.423 6.794 13.075 1.180 8.715